Black and Decker Corporation has the following information is for their standard and actual costs:

Standard Costs:

Budgeted units of production - 9,000 (100% of capacity)

Standard material per unit - 12 lbs.

Standard material cost - $14.00 per pound

Standard labor hours per unit - 12

Standard labor rate - $29.00 per hour

Standard variable overhead rate - $7.50 per labor hour

Budgeted fixed overhead - $729,000

Fixed overhead rate is based on budgeted labor hours at 100% capacity

Actual Costs:

Actual Production - 8,400 units

Actual material purchased and used - 87,360 pounds

Actual total material cost - $1,222,350

Actual labor - 101,200 hours

Actual total labor costs - $2,884,200

Actual fixed overhead - $756,250

Actual variable overhead - $715,000

Determine: (a) the quantity, price variance, and total direct materials cost variance; (b) the time variance, rate variance, and total direct labor cost variance; and (c) the volume variance, controllable variance, and total factory overhead cost variance.
Answer #1

Solution a:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
87360 $13.99 $1,222,350.00 87360 $14.00 $1,223,040.00 100800 $14.00 $1,411,200.00
$690.00 F $188,160.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $690.00 F
Direct material quantity variance $188,160.00 F
Direct material cost variance $188,850.00 F

Solution b:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
101200 $28.50 $2,884,200.00 101200 $29.00 $2,934,800.00 100800 $29.00 $2,923,200.00
$50,600.00 F $11,600.00 U
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $50,600.00 F
Direct Labor Efficiency variance $11,600.00 U
Direct labor cost variance $39,000.00 F

Solution c:

Controllable Variance
Actual overhead $1,471,250.00
Budgeted overhead $1,485,000.00
Controllable variance $13,750.00 F
Fixed overhead volume variance
Budgeted fixed overhead $729,000.00
Fixed overhead cost applied $680,400.00
Fixed overhead volume variance $48,600.00 U

Total factory overhead cost variance = $13,750 F + $48,600 U = $34,850 U